OpenStudy (anonymous):

2b+4=-18-9b Need some help Distributive Propertys i guess ?

OpenStudy (katrinakaif):

Your first step is to bring all of the constants to one side of the equal sign, and the variables to another.

OpenStudy (katrinakaif):

Lets first bring -9b to the other side by adding it with +2b 2b + 9b + 4 = 18 11b + 4 = 18 Solve like a linear equation 11b = 18-4 11b = 14 b = 14/11

OpenStudy (anonymous):

Thanks i am weak in this area lol

OpenStudy (katrinakaif):

No problemo! If you need additional help, or clarifications, simply post more questions! ;)

OpenStudy (anonymous):

I like this site it is new right ?

OpenStudy (katrinakaif):

Actually, its been up for more than a year or so. but Welcome to Openstudy! I found this site 8 months ago :P